OpenStudy (anonymous):

Solving systems using subtitution y=2x-10 2y=x-8

OpenStudy (anonymous):

First you plug the y=2x-10 into the y for 2y=x-8 This will give you 2(2x-10)=x-8 4x-20=x-8 3x=12 answer: x=4

OpenStudy (anonymous):

sub the first equation in to the second. 2(2x-10)=x-8 solve normally find x=4 then sub into first equation to find y

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

sure no prob

OpenStudy (anonymous):

i have to put it in (--) so would the x &y be (4,0) cause thts where i got confused

OpenStudy (anonymous):

ooooh. okay

OpenStudy (anonymous):

what grade are you in now? year 10?

OpenStudy (anonymous):

11 i have to retake algebra

OpenStudy (anonymous):

oooh.

OpenStudy (anonymous):

yea but im no good at math but my teacher helps me this just confuses me

OpenStudy (anonymous):

its okay. everyone has probs. i suck at chemistry

OpenStudy (anonymous):

I cant wait to get this course done

OpenStudy (anonymous):

lol. going of igcse? i have 2 years to go

OpenStudy (anonymous):

dang but would the x&y be put in the Parenthes's like that or did i miss something

OpenStudy (anonymous):

has to be put in parenthesis

OpenStudy (anonymous):

when you are writing coordinates, has to be in parenthesis

OpenStudy (anonymous):

i know that but im just wondering if y is zero so it would be (4,0)